# Tiling by Near Coincidence

### v1

@@ VI Summary and future directions
Inspired by physical bilayer systems and their moiré patterns, we have introduced a new method for generating quasiperiodic tilings of the plane. The method is both intuitive and rigorous: intuitive because it builds directly on the near coincidence of points from superimposed periodic layers, and rigorous because it can be mapped, in a variety of ways, to the established cut-and-project formalism. Using this framework we have reproduced several well-known quasiperiodic tilings, including the octagonal Ammann–Beenker, the dodecagonal Niizeki–Gähler, and the square and hexagonal Fibonacci tilings. Along the way, we have identified different ways in which the near-coincidence method corresponds to the cut-and-project approach: in some cases directly, in others after rescaling the coincidence window, and in the hexagonal Fibonacci case after recognizing the existence of several coincidence windows for distinct subsets of vertices. We note that other bilayer construction schemes were proposed over the years for generating quasiperiodic tilings, which bear some similarities with our near-coincidence method [Sadoc90, Stampfli12, Sadoc23].
@@
While the cut-and-project method remains far more powerful and general, the near-coincidence construction offers a perspective in which certain choices become natural that are rarely considered in the cut-and-project framework. Chief among these is the use of a circular coincidence window—the simplest and most intuitive choice in our method, but one that is not normally employed in cut-and-project constructions. In this way we have discovered new quasiperiodic tilings that would not likely have been encountered otherwise, like the tiling of Fig. 10d, containing the tristar tile  [^2] A similar tristar tiling is obtained by using a larger dodecagonal coincidence window. See the last row in the triangular bilayer section of Tab. 1.. Moreover, the circular window introduces an area proportional to $\pi$, and thus assigns transcendental values to the relative frequencies of local vertex configurations, in contrast to the algebraic frequencies associated with polygonal windows. This feature highlights an additional and unexpected richness of the method.
@@
An important observation concerns the diffraction spectrum, or equivalently the Fourier transform, of the near-coincidence tiling. In this framework, the bilayer diffracts as the superposition of the diffraction patterns of the rotated or rescaled individual layers, producing a mixed diffraction pattern that does not constitute a reciprocal lattice, or $\mathbb{Z}$-module, in the strict mathematical sense of closure under vector addition. However, the process of accepting only nearly coincident pairs of points and shifting them to their midpoints, which in physical systems may arise from interactions between the layers, generates additional harmonics at linear combinations of Bragg peaks from the separate layers. This gives rise to a well-behaved reciprocal lattice, exhibiting closure under vector addition along with the prescribed $N$-fold symmetry.
@@
A practical advantage of the near-coincidence method lies in its algorithmic simplicity. Once the layers are defined, the construction requires only the specification of a coincidence threshold and the testing of near coincidences. This lends itself naturally to numerical implementation. Indeed, we have developed an application, implemented by Ochana [Ochana25], that accepts as input the choice of layers, their twist or scaling, and the desired coincidence window—circular or polygonal—and generates the corresponding quasiperiodic tiling. This tool has already proved invaluable for exploring known examples and for discovering new ones. All parameters for the tilings discussed here are provided in Appendix A, and can be used directly within the application, presented in Appendix B.
@@ Acknowledgments
The authors are grateful to many colleagues, especially to Marianne Quiquandon and Denis Gratias, for their engaging comments during the $15^{th}$ International Conference on Quasicrystals (2023), where initial ideas about this work were first presented [Ochana23], and later during the conference on Algebra, Analysis, and Aperiodic Order (2024), dedicated to Michael Baake and Franz Gähler, on the occasion of their 65th birthdays. RL thanks Aviram Uri and Sergio de la Barrera for posing the challenge of assigning a proper quasiperiodic tiling to the trilayer graphene system, studied in Ref. [Uri23], which we intend to address in a future publication. This research is funded by the Israel Science Foundation (ISF), through grant number 1259/22.
Family Bilayer type Window parameters (shape, rotation) Edge length Threshold Figure
Octagonal (45 twist) 2 square lattices circle, 1.2, 2.23, 2.9 0.224 4a
octagon, 0, 2.91 0.224 4b
Dodecagonal (30 twist) 2 triangular lattices dodecagon, 15, 1.86 0.268 8b
dodecagon, 0, 0.299 8c
union of 2 hexagons, $\pm 15^{\circ}$, $\cup$ 0.299 8d
circle, 0.268 10b
circle, 0.299 10c
circle, 0.362 10d
dodecagon, 0, 0.374
3 square lattices dodecagon, 0, 2.15 0.29 15a
circle, 0.29
Fibonacci ($\tau$-scaled) 2 square lattices square, 0, 1.8, 3.11 0.437 11b
2 honeycomb tilings circle, 3.11, 4.92 0.383 14b
Table 1: Parameters for the quasiperiodic tilings generated by the near-coincidence method. The Table is grouped by family: octagonal tilings from 45 twisted bilayers, dodecagonal tilings from 30 twisted bilayers (and trilayers), and Fibonacci tilings from $\tau$-scaled bilayers. In each case we indicate the number and type of individual layers, the coincidence window (illustrated by a small symbol), the edge length, the threshold parameter, and a reference to the figure showing the tiling.
@@ Appendix B An interactive web application of the near-coincidence method
An interactive HTML demonstration of the near-coincidence method, developed with the assistance of the Gemini model [gemini23], is available at: https://meshyochana.github.io/near_coincidence/

### v2

@@ VI Summary and future directions
Inspired by physical bilayer systems and their moiré patterns, we have introduced a new method for generating quasiperiodic tilings of the plane. The method is both intuitive and rigorous: intuitive because it builds directly on the near coincidence of points from superimposed periodic layers, and rigorous because it can be mapped, in a variety of ways, to the established cut-and-project formalism. Using this framework we have reproduced several well-known quasiperiodic tilings, including the octagonal Ammann–Beenker, the dodecagonal Niizeki–Gähler, and the square and hexagonal Fibonacci tilings, the latter obtained by first analyzing the underlying one-dimensional construction. Along the way, we have identified different ways in which the near-coincidence method corresponds to the cut-and-project approach: in some cases directly, in others after rescaling the coincidence window, and in the hexagonal Fibonacci case after recognizing the existence of several coincidence windows associated with distinct vertex classes. We note that other bilayer construction schemes were proposed over the years for generating quasiperiodic tilings, which bear some similarities with our near-coincidence method [31, 32, 27].
@@
While the cut-and-project method remains far more powerful and general, the near-coincidence construction offers a perspective in which certain choices become natural that are rarely considered [25] in the cut-and-project framework. Chief among these is the use of a circular coincidence window—the simplest and most intuitive choice in our method, but one that is not normally employed in cut-and-project constructions. In this way we have discovered new quasiperiodic tilings that would not likely have been encountered otherwise, like the tiling of Fig. 10d, containing the tristar tile  [^3] A similar tristar tiling is obtained by using a larger dodecagonal coincidence window. See the last row in the triangular bilayer section of Tab. 1..
@@
An important observation concerns the diffraction spectrum, or equivalently the Fourier transform, of the near-coincidence tiling. In this framework, the bilayer diffracts as the superposition of the diffraction patterns of the rotated or rescaled individual layers, producing a mixed diffraction pattern that does not constitute a reciprocal lattice, or $\mathbb{Z}$-module, in the strict mathematical sense of closure under vector addition. However, the process of accepting only nearly coincident pairs of points and shifting them to their midpoints, which in physical systems may arise from interactions between the layers, generates additional harmonics at linear combinations of Bragg peaks from the separate layers. This gives rise to a well-behaved reciprocal lattice, exhibiting closure under vector addition along with the prescribed NN-fold symmetry.
@@
A practical advantage of the near-coincidence method lies in its algorithmic simplicity. Once the layers are defined, the construction requires only the specification of a coincidence threshold and the testing of near coincidences. This lends itself naturally to numerical implementation. Indeed, we have developed a web-based application, implemented by Ochana [34], that accepts as input the choice of layers, their twist or scaling, and the desired coincidence window—circular or polygonal—and generates the corresponding quasiperiodic tiling. This tool has already proved invaluable for exploring known examples and for discovering new ones. All parameters for the tilings discussed here are provided in Appendix A, and can be used directly within the application, presented in Appendix B.
@@ Acknowledgments
The authors are grateful to many colleagues, especially to Marianne Quiquandon and Denis Gratias, for their engaging comments during the $15^{th}$ International Conference on Quasicrystals (2023), where initial ideas about this work were first presented [35], and later during the conference on Algebra, Analysis, and Aperiodic Order (2024), dedicated to Michael Baake and Franz Gähler, on the occasion of their 65th birthdays. RL thanks Aviram Uri and Sergio de la Barrera for posing the challenge of assigning a proper quasiperiodic tiling to the trilayer graphene system, studied in Ref. [16], which we intend to address in a future publication. This research is funded by the Israel Science Foundation (ISF), through grant number 1259/22.
@@ Appendix B An interactive web application of the near-coincidence method
An interactive HTML demonstration of the near-coincidence method, developed with the assistance of the Gemini model [40], is available at: https://meshyochana.github.io/near_coincidence/